# HarmonyCut: Supporting Creative Chinese Paper-cutting Design with Form and Connotation Harmony

### v1

@@ 3.1. Participants
As a work targeted to public users, we initially sought to examine the paper-cutting design process and the challenges encountered without GenAI. Subsequently, we aimed to assess the role and limitations of integrating GenAI into paper-cutting design. To achieve this, we recruited participants with varying levels of expertise in both paper-cutting and GenAI, as experts are expected to have insights into design workflows, and each participant group faces different challenges in paper-cutting design. We categorized participants into four levels of paper-cutting expertise: (1) Masters: who have over 20 years of professional experience in paper-cutting creation are officially recognized as ICH inheritors; (2) Practitioners: who have 10-20 years of experience in paper-cutting-related work; (3) Amateurs: who have 1-3 years of experience in creating paper-cuttings without systematic training; (4) Novices: who never engaged in paper-cutting design or creation. Additionally, we defined three levels of GenAI expertise: (1) Professionals: researchers in the multi-modal machine learning field; (2) Knowledgeable Users: who regularly integrate GenAI into their professional, educational, or personal activities and have some experience with basic prompt engineering. (3) Novices: who have minimal exposure to GenAI, may have heard of it without engaging in its use, or are entirely unacquainted with it.
@@ 4. Content Analysis of Paper-cuttings and Patterns
Based on the formative study, we discovered that there are some key aspects to consider in the design workflow. To identify them, we first collected a paper-cutting dataset. We conducted two content analyses: one to develop an ideation factors taxonomy and another for a pattern taxonomy (taxonomy for the key element in ideation and composition). Both analyses were carried out under the guidance of expert review, involving five experts (P1–P5) who had also participated in the formative study. The experts were involved in two stages: (1) During the coding process, ambiguities in the instance-level annotation of paper-cuttings and patterns were resolved through a single expert-guided discussion conducted via the WeChat group, with the final annotation determined based on the majority vote of all experts; (2) Additionally, after each version of the codebook was drafted by the authors, the experts participated in discussions conducted via online conferences. In each round, the experts collaboratively reviewed and evaluated the current version of the codebook, offering refined or expanded suggestions, which were systematically discussed and consolidated to resolve ambiguities and ensure alignment of perspectives. This iterative process continued until all experts reached a consensus on the coding results, ensuring that the final codebook was validated and more comprehensive.
@@ 5. Design Goals
    Reference-based Exploration for Composition For each compositional action, the system suggests related content, including both composite and unit patterns in paper-cutting. Users can explore these design references for creative self-selection and arrangement of contours and patterns (C3, C4).
@@
    Controllable and Editable Design with Recommendation The recommendations provided by GenAI are integral to DG 1 and 2. These suggestions must ensure diversity to foster creativity while maintaining rationality and relevance, thereby preventing users from being misled or overwhelmed during exploration (C4). Furthermore, users are able to edit selections of unreasonable or partially reasonable ideas and compositions, allowing them to incorporate desired elements into their designs and complete the entire process (C5).
6. HarmonyCut
@@ 6.4. Implementation Details
Our system utilizes a Flask [^8] https://flask.palletsprojects.com-based back-end, integrated with the GLM-4 APIs for text processing and DALL-E-3 for paper-cutting image generation. The front-end is built using Vite + Vue, with additional support from Element Plus for UI components, Pinia for state management, and Tailwind CSS for styling. A fine-tuned CLIP model for multi-modal retrieval, ViT-base-16 and RoBERTa-wwm-base are used for visual and text encoding. For image segmentation, we extended the segment-anything-web UI [^9] https://github.com/Kingfish404/segment-anything-webui framework, which enables extraction and manipulation of image components. The backend supports the use of SAM-ViT-base [^10] https://github.com/facebookresearch/segment-anything (Kirillov et al., 2023) by default, with flexible configurations for alternative models. Additionally, a fine-tuned pre-trained ViT (ViT-base-16, pre-trained on ImageNet-21k) is employed for unit-pattern classification. The frontend supports dynamic SVG manipulation and bitmap-to-SVG conversion using Fabric [^11] https://fabricjs.com/ and Potrace [^12] https://potrace.sourceforge.net/ libraries. We also implemented custom features such as object dragger and pattern editor, integrated with the overall canvas design interface. The related dataset, comprising 4,031 text-image pairs for fine-tuning the multi-modal retrieval model and 1,279 annotated images for fine-tuning the unit pattern recognition model, is detailed in the supplementary material.
Table 3. Prepared questions during expert interviews from 3 topics.
[…]
@@ 7.2. Procedure and Measures
To validate the design workflow integrating paper-cutting knowledge with HarmonyCut, participants completed a questionnaire after each session, resulting in two submissions per participant. The questionnaire included three 6-point Likert scale questions focused on design goals: (1) ideation guidance, asking if the tool helps generate ideas; (2) exploration with reference, asking if the tool supports broad exploration for design, complementary to the Exploration measure in CSI; and (3) editing flexibility, questioning if the tool allows sufficient editing flexibility. Additionally, the questionnaire incorporated the CSI (Cherry and Latulipe, 2014) and NASA-TLX (Hart, 2006) to evaluate usability regarding creative support and perceived workload. Each CSI metric is rated on a scale of 0-10, while each NASA-TLX metric is rated on a scale of 0-20.
Prior to the expert interview, each expert was asked to prepare a design intent. During the interview, experts received a 15-minute introduction and system walk-through, including a specific design task. After learning how to use HarmonyCut for the paper-cutting design task, experts were given 30 minutes to create a paper-cutting based on their prepared design intent. Following the design session, an interview focused on three topics adapted from Xu et al. (Xu et al., 2023), with each topic consisting of two questions to collect their feedback.
@@ 8.2. Integrate the Cultural Knowledge with GenAI Support Creative Design
GenAIs have demonstrated strong capabilities in both language and visual understanding, allowing users to engage in the creative process through natural language input. However, in tasks involving abstract cultural and artistic design, these models often fail to fully capture user intent, resulting in outputs that may not meet expectations. The findings from this study, alongside insights from previous research (Messer, 2024; Garcia, 2024; Chung and Adar, 2023a), highlight the limitations of GenAI in comprehending and interpreting cultural elements. To prevent these shortcomings from impeding creative support, it is essential to incorporate cultural knowledge into the models. Although large-scale data augmentation offers one possible solution, it is limited by both the scale and complexity of data collection. E1 emphasized that the taxonomy used to organize knowledge through design factors effectively captures a significant portion of paper-cutting knowledge. However, the vast and intricate nature of cultural knowledge makes it impractical to rely solely on dataset expansion to improve GenAI’s ability to support culturally relevant design. A more effective approach may involve structuring cultural content. Previous work in the field of painting (Chung and Adar, 2023b) has attempted to extract vague concepts and process them individually, but these models often lack the abstract knowledge necessary, resulting in a disconnect between abstract ideas and concrete visual outputs. By structuring abstract content into design tasks and design spaces, models can generate more appropriate outputs for cultural art-related tasks. In this study, HarmonyCut organizes the paper-cutting design process around factors and patterns, while Magical Brush uses symbols as fundamental elements in Chinese painting (Xu et al., 2023). Similarly, template-based methods have been applied to dynamic shadow puppetry creation (Yao et al., 2024). These structured approaches provide a foundational framework for complex and profound cultural knowledge, allowing models to better understand with cultural content, such as the visual selection and composition in HarmonyCut, partially bridging gaps in expertise and cultural background. This makes it possible for public users to access relevant knowledge, participate in cultural creative design, and even further develop expertise, as supported by the feedback from E1–E3 agreeing with the cultural aspects in Table 3, thereby engaging and supporting cultural communication while sustaining the vitality of cultural content such as ICH.
@@ B.2.1. For the Codebook of Core Factors of Paper-cutting Design Ideation
The initial version of the codebook was developed solely based on style and function. After the first round of discussions, the experts recommended adding two dimensions, including subject matter (P1-P2 and P4-P5) and method of expression (P2 and P4-P5), which were agreed upon through consensus. During the second round, they proposed consolidating categories such as plants, animals, and mythical creatures into a broader “flora and fauna” type. By the third round of discussions, the experts collectively agreed that the current version of the codebook was generally reasonable.
@@ B.2.2. For the Codebook of Patterns in Paper-cutting
During the first round of discussions, the experts noted that, beyond unit geometry and sawtooth patterns, many Unit Patterns, such as crescent, cloud, and fire patterns, have more practical semantics rather than abstract ones. Therefore, they recommended adding a new sub-category under Unit Patterns: unit semantic patterns. For Composite Patterns, the experts proposed distinguishing a type focused on functionality and expressing core themes, referred to as composite primary patterns, in contrast to decorative patterns. After the second round of discussions, they concluded that the current version of the codebook was generally reasonable.
@@ B.5. Coding Results of the Patterns of Paper-cuttings
This figure shows coding distribution results The coding distribution results: 635 selections and 1269 cut-outs for the Unit Pattern, 70 selections and 140 paper-cuttings for the Composite Pattern, which validate the consistency of coding from the construction to the result.

### v2

@@ 3.1. Participants
As a work targeted to public users, we initially sought to examine the paper-cutting design process and the challenges encountered without GenAI. Subsequently, we aimed to assess the role and limitations of integrating GenAI into paper-cutting design. To achieve this, we recruited participants with varying levels of expertise in both paper-cutting and GenAI, as experts are expected to have insights into design workflows, and each participant group faces different challenges in paper-cutting design. Participants were categorized into four levels of paper-cutting expertise, with further clarification regarding the criteria for paper-cutting expertise provided in subsection A.1: (1) Masters: who have over 20 years of professional experience in paper-cutting creation are officially recognized as ICH inheritors; (2) Practitioners: who have 10-20 years of experience in paper-cutting-related work; (3) Amateurs: who have 1-3 years of experience in creating paper-cuttings without systematic training; (4) Novices: who never engaged in paper-cutting design or creation. Additionally, we defined three levels of GenAI expertise: (1) Professionals: researchers in the multi-modal machine learning field; (2) Knowledgeable Users: who regularly integrate GenAI into their professional, educational, or personal activities and have experience with basic prompt engineering. (3) Novices: who have minimal exposure to GenAI, may have heard of it without engaging in its use, or are entirely unacquainted with it.
@@ 4. Content Analysis of Paper-cuttings and Patterns
Based on the formative study, we discovered that there are some key aspects to consider in the design workflow. To identify them, we first collected a paper-cutting dataset. We conducted two content analyses: one to develop an ideation factors taxonomy and another for a pattern taxonomy (taxonomy for the key element in ideation and composition). Both analyses were carried out under the guidance of expert review, involving five experts (P1–P5) who had also participated in the formative study. The experts were involved in two stages: (1) During the coding process, ambiguities in the instance-level annotation of paper-cuttings and patterns were resolved through a single expert-guided discussion conducted via the WeChat group, with the final annotation determined based on the majority vote of all experts; (2) Additionally, after each version of the codebook was drafted by the authors, the experts participated in discussions conducted via online conferences. In each round, the experts collaboratively reviewed and evaluated the current version of the codebook, offering refined or expanded suggestions, which were systematically discussed and consolidated to resolve ambiguities and ensure alignment of perspectives. This iterative process continued until all experts reached a consensus on the coding results, ensuring that the final codebook was validated.
@@ 5. Design Goals
    Reference-based Exploration for Composition For each compositional action, the system based on domain knowledge suggests related content as the reference. Users can explore these design references for creative self-selection and arrangement of contours and patterns extracted from reference (C3, C4).
@@
    Controllable and Editable Design with Recommendation The recommendations provided by GenAI are integral to DG 1 and 2. These suggestions can ensure diversity to foster creativity while maintaining rationality and relevance, thereby preventing users from being misled or overwhelmed during exploration (C4). Furthermore, users are able to edit selections of unreasonable or partially reasonable ideas and compositions, allowing them to incorporate desired elements into their designs and complete the entire design process (C5).
@@ 6.4. Implementation Details
Our system utilizes a Flask [^8] https://flask.palletsprojects.com-based back-end, integrated with the GLM-4 APIs for text processing and DALL-E-3 for paper-cutting image generation. The front-end is built using Vite + Vue, with additional support from Element Plus for UI components, Pinia for state management, and Tailwind CSS for styling. A fine-tuned CLIP model for multi-modal retrieval, ViT-base-16 and RoBERTa-wwm-base are used for visual and text encoding. For image segmentation, we extended the segment-anything-web UI [^9] https://github.com/Kingfish404/segment-anything-webui framework, which enables extraction and manipulation of image components. The backend supports the use of SAM-ViT-base [^10] https://github.com/facebookresearch/segment-anything (Kirillov et al., 2023) by default, with flexible configurations for alternative models. Additionally, a fine-tuned pre-trained ViT (ViT-base-16, pre-trained on ImageNet-21k) is employed for unit-pattern classification. The frontend supports dynamic SVG manipulation and bitmap-to-SVG conversion using Fabric [^11] https://fabricjs.com/ and Potrace [^12] https://potrace.sourceforge.net/ libraries. We also implemented custom features such as object dragger and pattern editor, integrated with the overall canvas design interface. The related datasets, comprising 140 paper-cutting images, 4,031 text-image pairs for fine-tuning the multi-modal retrieval model and 1,279 annotated images for fine-tuning the unit pattern recognition model, are detailed in the supplementary material.
@@ 7.2. Procedure and Measures
To validate the design workflow integrating paper-cutting knowledge with HarmonyCut, participants completed a questionnaire after each session, resulting in two submissions per participant. The questionnaire included three 6-point Likert scale questions focused on design goals: (1) ideation guidance, asking if the tool helps generate ideas; (2) exploration with reference, asking if the tool supports broad exploration for design, complementary to the Exploration measure in CSI; and (3) editing flexibility, questioning if the tool allows sufficient editing flexibility. Additionally, the questionnaire incorporated the CSI (Cherry and Latulipe, 2014) (0-10 scale) and NASA-TLX (Hart, 2006) (0-20 scale) to evaluate usability regarding creative support and perceived workload.
Prior to the expert interview, each expert was asked to prepare a design intent. During the interview, experts received a 15-minute introduction and system walk-through, including a specific design task. After learning how to use HarmonyCut for the paper-cutting design task, experts were given 30 minutes to create a paper-cutting based on their prepared design intent. After the design session, an interview was conducted on three topics adapted from Xu et al.(Xu et al., 2023) (Table 3), each with two questions to gather feedback.
Table 3. Questions during expert interviews from 3 topics.
[…]
@@ 8.2. Integrate the Cultural Knowledge with GenAI Support Creative Design
GenAIs have demonstrated strong capabilities in both language and visual understanding, allowing users to engage in the creative process through natural language input. However, in tasks involving abstract cultural and artistic design, these models often fail to fully capture user intent, resulting in outputs that may not meet expectations. The findings from this study, alongside insights from previous research (Messer, 2024; Garcia, 2024; Chung and Adar, 2023a), highlight the limitations of GenAI in comprehending and interpreting cultural elements. To prevent these shortcomings from impeding creative support, it is essential to incorporate cultural knowledge into the models. Although large-scale data augmentation offers one possible solution, it is limited by both the scale and complexity of data collection. E1 emphasized that the taxonomy used to organize knowledge through design factors effectively captures a significant portion of paper-cutting knowledge.
However, the vast and intricate nature of cultural knowledge makes it impractical to rely solely on dataset expansion to improve GenAI’s ability to support culturally relevant design. A more effective approach may involve structuring cultural content. Previous work in the field of painting (Chung and Adar, 2023b) has attempted to extract vague concepts and process them individually, but these models often lack the abstract knowledge necessary, resulting in a disconnect between abstract ideas and concrete visual outputs. By structuring abstract content into design tasks and design spaces, models can generate more appropriate outputs for cultural art-related tasks.
In this study, HarmonyCut organizes the paper-cutting design process around factors and patterns, while Magical Brush uses symbols as fundamental elements in Chinese painting (Xu et al., 2023). Similarly, template-based methods have been applied to dynamic shadow puppetry creation (Yao et al., 2024). These structured approaches provide a foundational framework for complex and profound cultural knowledge, allowing models to better understand with cultural content, such as the visual selection and composition in HarmonyCut, partially bridging gaps in expertise and cultural background. This makes it possible for public users to access relevant knowledge, participate in cultural creative design, and even further develop expertise, as supported by the feedback from E1–E3 agreeing with the cultural aspects in Table 3, thereby engaging and supporting cultural communication while sustaining the vitality of cultural content such as ICH. While our structured approach, which relies on manual annotation and model fine-tuning, enhances the model’s ability to understand cultural knowledge, it remains limited in achieving comprehensive coverage of large-scale knowledge, as further discussed in subsection 8.3.
@@ Appendix A Formative Study
A.1. Expertise Levels of Paper-cutting
As mentioned in Appendix A, the key criterion for distinguishing experts (i.e., master and practitioner) is whether they have undergone systematic training (specifically, apprenticing under an inheritor) or have transitioned from amateur to experts status through the accumulation of substantial paper-cutting experience and even been finally recognized as an inheritor. In our interviews, P1-P4 highlighted that they developed an interest in paper-cutting as early as 3-4 years old. Their families subsequently petitioned inheritors to accept these children as apprentices, and the acceptance of such requests indicated the masters’ recognition of their talent. As a result, they began receiving systematic and professional training by the age of 5-9. This practice of training young children is relatively common in the Chinese paper-cutting inheritance. However, because this tradition relies on early engagement, the decreasing interest among the younger generation has partially contributed to the challenges faced in the inheritance of ICH, including paper-cutting.
@@ B.2.1. For the Codebook of Core Factors of Paper-cutting Design Ideation
The initial version of the codebook was developed solely based on style and function. Thus, in the first round of discussions, the experts recommended adding two dimensions, including subject matter (P1-P2 and P4-P5) and method of expression (P2 and P4-P5), which were accepted through all experts’ consensus. During the second round, P2 and P5 suggested merging specific types, such as plants, animals, and mythical creatures into a broader “flora and fauna” type. By the third round of discussions, the experts collectively agreed that the current version of the codebook was generally adequate.
@@ B.2.2. For the Codebook of Patterns in Paper-cutting
During the first round of discussions, the experts observed that, in addition to unit geometry and sawtooth patterns, many Unit Patterns, such as crescent, cloud, and fire patterns, possess more practical semantics rather than purely abstract ones. As a result, they recommended introducing a new sub-category within Unit Patterns, termed unit semantic patterns. Regarding Composite Patterns, the experts suggested distinguishing a type that emphasizes functionality and conveys core themes, referred to as composite primary patterns, as opposed to decorative patterns. Following the second round of discussions, the experts collectively agreed that the current version of the codebook was generally appropriate.
@@ B.5. Coding Results of the Patterns of Paper-cuttings
This figure shows coding distribution results. The coding distribution results: 635 selections and 1269 cut-outs for the Unit Pattern, 70 selections and 140 paper-cuttings for the Composite Pattern, which validate the consistency of coding from the construction to the result.